PLEASE SHOW STEPS, I have the answers, I just need a step by step guide

4. Having laser beams in the place of eyes is an x-linked recessive condition. In a (made up) population, 35% of the males has the laser beam condition (Xb).

a. What is the allele frequency in this population?

b. What is the probability that a random male and random female in the population will have a girl with the laser beam condition?
Answer #1

Answer:

I4a). In X-linked recessive inheritance, the frequency of trait in male is equal to frequency of allele as males have only one X-chromosome.

Laser beam allele frequency = q^2=35% = 0.35

b).

Frequency of laser beam in female =q^= 0.35 * 0.35 = 0.1225

The frequency of girl child is 1/2 or 0.5 as two genders (male or female) and each gender has equal proportion.

The probability of girl with the laser beam condition = 0.35 * 0.35 * 0.5 = 0.06125 or 6.125%